Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 274     Entry time: 25 Jul 2023, 19:18
Author: Harald Sinn 
Group: PRC 
Subject: Status 

After the 12 hour cryo failure last night, it was decided to exchange a faulty window with some reduced transmission at the injector laser today (see email by Andreas to xo-info yesterday). 

This was sucessful and the injector tunnel was closed today 16:40 and restart of the gun is ongoing. Tomorrow ramp up and SASE tuning.

The hope is that the large jitter and the poor SASE performance of last week can be now improved.  Beam delivery to instrument is expected on Thursday. 

ELOG V3.1.4-7c3fd00